OpenStudy (anonymous):

Let f(x) = x – 2 and g(x) = 3x + 4. What is f(g(x)) ?

OpenStudy (ghazi):

to find f(g(x)) put x= g(x) in f(x) can you do it?

OpenStudy (anonymous):

a little confused

OpenStudy (ghazi):

well put x=3x+4 in f(x)

OpenStudy (ghazi):

f(x)= (3x+4)-2

OpenStudy (ghazi):

@marci123 is it clear now?

OpenStudy (across):

\[ f(x)=x-2\\ f(g(x))=g(x)-2\\ f(g(x))=3x+4-2\\ f(g(x))=3x+2 \]

OpenStudy (anonymous):

thanks!!

OpenStudy (ghazi):

:)